OpenStudy (anonymous):

x^2-2x+1 what numbers make y=4

OpenStudy (lalaly):

3

OpenStudy (anonymous):

and what else ?

OpenStudy (lalaly):

x^2-2x+1 = 4 x^2 -2x -3=0 (x-3)(x+1) x=3 , x=-1